# Statistical Collusion by Collectives on Learning Platforms

### v1

@@ Appendix D Lemmas
Lemma D.2 (Hardt et al. (2023), Lemma B.1).
@@ Appendix E Proofs and additional remarks
The proof of Theorem 3.5 is essentially the same than the proof of Theorem 3.3. One difference is that we can not simplify $\underset{x\sim\mathcal{D}}{\mathbb{P}}(\hat{f}(g(x))=y^{*})=\underset{\tilde{x}\sim\tilde{\mathcal{D}}}{\mathbb{P}}(\hat{f}(\tilde{x})=y^{*})$ as we did in the proof of Theorem 3.3. Therefore, the calculations must be carried out directly using the term $g(x)$.
@@
by applying Hoeffding’s inequality to both empirical probabilities. Note that Hoeffding’s inequality for the term $\underset{(x,y)\sim\mathcal{D}}{\mathbb{P}}(x=\tilde{x},y=\hat{y}_{\tilde{x}})$ holds in probability over $D^{(N-n)}$ conditionnally on $D^{(n_{\rm est})}$. It should only be applied #$\#\tilde{\mathcal{X}}$ times here contrary to #$\#\tilde{\mathcal{X}}(\#\mathcal{Y}-1)$ times in signal planting because $\hat{y}_{\tilde{x}}$ is fixed conditionnally on $D^{(n_{\rm est})}$. Then, since the inequality holds in probability over $D^{(N-n)}$ conditionnally on $D^{(n_{\rm est})}$, in particular it holds marginally over all consumers.
The conclusion is similar to that of Theorem 3.3. We just need to apply the same argument as above to use Hoeffding’s inequality: $\underset{(x,y)\sim\mathcal{D}}{\mathbb{P}}(x=\tilde{x},y=\hat{y}_{\tilde{x}})\geq\underset{(x,y)\sim D^{(n-n_{\rm est})}}{\hat{\mathbb{P}}}(x=\tilde{x},y=\hat{y}_{\tilde{x}})-R_{\tilde{\delta}}(n-n_{\rm est})$. ∎
In fact, it is possible to obtain a slightly better bound than (4). Indeed, to bound $\Delta_{\tilde{x}}^{(\mathcal{D})}$, one can use Hoeffding’s inequality on $D^{(n)}$ instead of $D^{(n-n_{\rm est})}$ for the term $\underset{(x,y)\sim\mathcal{D}}{\mathbb{P}}(x=\tilde{x},y=y^{*})$. We kept the inequality (4) as it is to simplify the expression of the bound and make the interpretation of the different terms easier.
@@
  • $-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n-n_{\rm est})}+2R_{\tilde{\delta}}(n-n_{\rm est})+2R_{\tilde{\delta}}(N-n)\right)$
@@
The first and third terms play the same role as in signal planting. The main difference here is the second term $-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n-n_{\rm est})}+2R_{\tilde{\delta}}(n-n_{\rm est})+2R_{\tilde{\delta}}(N-n)\right)$. The term $\Delta_{\tilde{x}}^{(n-n_{\rm est})}$ represents how much more probable the target label $y^{*}$ is compared to other labels given a feature belonging to $\tilde{\mathcal{X}}$. If $y^{*}$ is much more probable than the other labels, it will be difficult to unplant the signal. Conversely, if $y^{*}$ is not the most probable, it will be easier to unplant the signal. The collective’s plan relies on the choice of $n_{\rm est}<n$. If $n_{\rm est}$ is large, it will be more likely for the collective to select the best possible label given a modified feature $\tilde{x}$ to maximize the impact of the strategy $h$. However, if $n_{\rm est}$ is too large compared to $n$, the estimation error $2R_{\tilde{\delta}}(n-n_{\rm est})$ may become too significant. Therefore, there is a trade-off for the optimal choice of the value $n_{\rm est}$. This trade-off depends on the specific dataset being considered.
@@
$\hat{S}(n)\geq\underset{x^{\prime}\sim\mathcal{D}}{\mathbb{P}}\left(\frac{n}{N}\left(\underset{x\sim\mathcal{D}}{\mathbb{P}}(x=x^{\prime})-R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{x^{\prime}}^{(\mathcal{D})}+2R_{\tilde{\delta}}(N-n)\right)>0\right)-R_{\tilde{\delta}}(N_{\rm test})$
@@
$\hat{S}(n)\geq\underset{x^{\prime}\sim D^{(n)}}{\mathbb{P}}\left(\frac{n}{N}\left(\underset{x\sim\mathcal{D}}{\mathbb{P}}(x=x^{\prime})-R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{x^{\prime}}^{(\mathcal{D})}+2R_{\tilde{\delta}}(N-n)\right)>0\right)-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}).$
@@ Appendix F Bounds in the Infinite Data Regime
In the infinite data regime, the framework is the same than in algorithmic collective action: the platform observes the distribution
@@ F.2 Signal unplanting
The signal unplanting strategy relies on the choice of some parameter $n_{\rm est}<n$ to estimate the most likely label $y_{g(x)}\neq y^{*}$ given some feature $g(x)$. In the infinite data regime, we assume that $n_{\rm est}\rightarrow\infty$ and $n-n_{\rm est}\rightarrow\infty$. The bound (4) from Theorem 3.7 becomes:

### v2

@@ Appendix D Lemmas
Lemma D.2 ((Hardt et al., 2023), Lemma B.1).
@@ Appendix E Proofs and additional remarks
The proof of Theorem 3.5 is essentially the same as the proof of Theorem 3.3. One difference is that we can not simplify $\underset{x\sim\mathcal{D}}{\mathbb{P}}(\hat{f}(g(x))=y^{*})=\underset{\tilde{x}\sim\tilde{\mathcal{D}}}{\mathbb{P}}(\hat{f}(\tilde{x})=y^{*})$ as we did in the proof of Theorem 3.3. Therefore, the calculations must be carried out directly using the term $g(x)$.
@@
by applying Hoeffding’s inequality to both empirical probabilities. Note that Hoeffding’s inequality for the term $\underset{(x,y)\sim\mathcal{D}}{\mathbb{P}}(x=\tilde{x},y=\hat{y}_{\tilde{x}})$ holds in probability over $D^{(N-n)}$ conditionally on $D^{(n_{\rm e})}$. It should only be applied #$\#\tilde{\mathcal{X}}$ times here contrary to #$\#\tilde{\mathcal{X}}(\#\mathcal{Y}-1)$ times in signal planting because $\hat{y}_{\tilde{x}}$ is fixed conditionally on $D^{(n_{\rm e})}$. Then, since the inequality holds in probability over $D^{(N-n)}$ conditionally on $D^{(n_{\rm e})}$, in particular it holds marginally over all consumers.
The conclusion is similar to that of Theorem 3.3. We just need to apply the same argument as above to use Hoeffding’s inequality: $\underset{(x,y)\sim\mathcal{D}}{\mathbb{P}}(x=\tilde{x},y=\hat{y}_{\tilde{x}})\geq\underset{(x,y)\sim D^{(n-n_{\rm e})}}{\hat{\mathbb{P}}}(x=\tilde{x},y=\hat{y}_{\tilde{x}})-R_{\tilde{\delta}}(n-n_{\rm e})$. ∎
In fact, it is possible to obtain a slightly better bound than (4). Indeed, to bound $\Delta_{\tilde{x}}^{(\mathcal{D})}$, one can use Hoeffding’s inequality on $D^{(n)}$ instead of $D^{(n-n_{\rm e})}$ for the term $\underset{(x,y)\sim\mathcal{D}}{\mathbb{P}}(x=\tilde{x},y=y^{*})$. We kept the inequality (4) as it is to simplify the expression of the bound and make the interpretation of the different terms easier.
@@
  • $-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n-n_{\rm e})}+2R_{\tilde{\delta}}(n-n_{\rm e})+2R_{\tilde{\delta}}(N-n)\right)$
@@
The first and third terms play the same role as in signal planting. The main difference here is the second term $-\frac{N-n}{N}\left(\Delta_{\tilde{x}}^{(n-n_{\rm e})}+2R_{\tilde{\delta}}(n-n_{\rm e})+2R_{\tilde{\delta}}(N-n)\right)$. The term $\Delta_{\tilde{x}}^{(n-n_{\rm e})}$ represents how much more probable the target label $y^{*}$ is compared to other labels given a feature belonging to $\tilde{\mathcal{X}}$. If $y^{*}$ is much more probable than the other labels, it will be difficult to unplant the signal. Conversely, if $y^{*}$ is not the most probable, it will be easier to unplant the signal. The collective’s plan relies on the choice of $n_{\rm e}<n$. If $n_{\rm e}$ is large, it will be more likely for the collective to select the best possible label given a modified feature $\tilde{x}$ to maximize the impact of the strategy $h$. However, if $n_{\rm e}$ is too large compared to $n$, the estimation error $2R_{\tilde{\delta}}(n-n_{\rm e})$ may become too significant. Therefore, there is a trade-off for the optimal choice of the value $n_{\rm e}$. This trade-off depends on the specific dataset being considered.
@@
$\begin{split}&\hat{S}(n)\geq\underset{x^{\prime}\sim\mathcal{D}}{\mathbb{P}}\left(\frac{n}{N}\left(\underset{x\sim\mathcal{D}}{\mathbb{P}}(x=x^{\prime})-R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{x^{\prime}}^{(\mathcal{D})}+2R_{\tilde{\delta}}(N-n)\right)>0\right)\\ &-R_{\tilde{\delta}}(N_{\rm test})\end{split}$
@@
$\begin{split}&\hat{S}(n)\geq\underset{x^{\prime}\sim D^{(n)}}{\mathbb{P}}\left(\frac{n}{N}\left(\underset{x\sim\mathcal{D}}{\mathbb{P}}(x=x^{\prime})-R_{\tilde{\delta}}(n)\right)-\frac{N-n}{N}\left(\Delta_{x^{\prime}}^{(\mathcal{D})}+2R_{\tilde{\delta}}(N-n)\right)>0\right)\\ &-R_{\tilde{\delta}}(n)-R_{\tilde{\delta}}(N_{\rm test}).\end{split}$
@@ Appendix F Bounds in the Infinite Data Regime
In the infinite data regime, the framework is the same as in algorithmic collective action: the platform observes the distribution
@@ F.2 Signal unplanting
The signal unplanting strategy relies on the choice of some parameter $n_{\rm e}<n$ to estimate the most likely label $y_{g(x)}\neq y^{*}$ given some feature $g(x)$. In the infinite data regime, we assume that $n_{\rm e}\rightarrow\infty$ and $n-n_{\rm e}\rightarrow\infty$. The bound (4) from Theorem 3.7 becomes: